TRIANGE ABC IF ANGLE A =X,ANGLE B=3x-2AND ANGLE C=Y AND ANGLE C- ANGLE B=9.FIND THREE ANGLES

Grade:10

1 Answers

Vijay Mukati
8 years ago
Hint, Using the conditionANGLE C- ANGLE B=9, we can find that angle C = 7 + 3x. So the three angles in terms of x will be A = x and B = 3x-2 and C = 7 + 3x. Thanks